OpenStudy (anonymous):

Given that X~ B(20,0.14), calculate: P(X

OpenStudy (anonymous):

\[P(X\le\ E(X)).\]

OpenStudy (zarkon):

\[\mathbb{E}(X)=np=20\times.14=2.8\] \[P(X\le\ \mathbb{E}(X))=P(X\le 2.8)=P(X\le2)=\sum_{x=0}^{2}{20\choose x}(.14)^x(1-.14)^{20-x}\approx.455\]